# Towards Domain Adaptive Neural Contextual Bandits

### v1

@@ A.2 Proof for the Target Regret Bound
$\mathit{R}_{\mathcal{T}}$ $\triangleq\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\Bigg{)}$
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are defined in Sec. 3. $\psi$ and $C$ are constants.
@@
[Please check the full proof in the next page.]
@@
$\mathit{R}_{\mathcal{T}}$ $=\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$=\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\epsilon_{\mathcal{T}}(h)+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
@@
$\leq\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
@@
$=\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$

### v2

@@ A.1 Proofs for Lemmas
Lemma A.4.
Denoting the contexts from the target domain ${\mathcal{T}}$ as $\{{\bf x}_{t,a}^{\mathcal{T}}\}_{i\in[N],a\in[K]}$ and the associated ground-truth action as $\{a_{i}^{*}\}_{i=1}^{N}$, the estimated regret (i.e., using the trained model $\widehat{\theta}$ from the source domain) for the target domain is
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)$ $\leq\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{T}},$
where $\kappa_{i}^{\mathcal{T}}$ is $\max\left\{2\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}+\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right\}$.
Proof.
By definition, we have
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)$
$=\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}-\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}+\alpha\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}-\alpha\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right|\right)$
$=\sum^{N}_{i=1}\left(\left|\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)-\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)\right|\right)$
$\quad+\sum^{N}_{i=1}\alpha\left(\left|\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}-\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right|\right)$
$\leq\sum^{N}_{i=1}\left(\left|\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)-\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)\right|\right)$
$\quad+\sum^{N}_{i=1}\alpha\max\left\{\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right\}$ (13)
$\leq\sum^{N}_{i=1}\left(\left|\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)\right|\right)+\sum^{N}_{i=1}\alpha\max\left\{\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right\}$ (14)
$\leq\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\max\left\{2\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}+\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right\}$
:-$\coloneq\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{T}},:$ (15)
where the inequality in Eqn. (13) is based on the fact that for any $a,b>0$, $\left|a-b\right|\leq\max\{a,b\}$. The inequality in Eqn. (14) is derived from the definition of $\widehat{a}_{i}$, since $\widehat{a}_{i}$ is selected from the maximum value of $\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,a}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)$, therefore we guarantee that $\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)\geq\left(\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle+\alpha\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right)$. In addition,  Eqn. (15) holds by definition as we define $\kappa_{i}^{\mathcal{T}}$ as $\max\left\{2\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}+\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\right\|_{A_{i-1}^{-1}}\right\}$. ∎
Next, we discuss the property of $\kappa_{i}^{\mathcal{T}}$ in the following lemma.
Lemma A.5.
For any arbitrary domain $\mathcal{D}$, we denote
$\kappa_{i}^{\mathcal{D}}=\max\left\{2\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}}+\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}}\right\}$. Then, by our DABand algorithm, as the time step $i\to\infty$, we have $\kappa_{i}^{\mathcal{D}}\to 0$.
Proof.
By definition of $A$, its singular value increase as $i$ increase, which means that the singular values in its inverse matrix, $A^{-1}$, will decrease to 00. Therefore, for any new array $u$, $||u||_{A^{-1}}=\sqrt{uA^{-1}u^{T}}\to 0$. Then when $i\to\infty$, we have
$\lim_{i\to\infty}\kappa_{i}^{\mathcal{D}}$ $=\max\left\{2\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}},\left\|\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}}+\left\|\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\right\|_{A_{i-1}^{-1}}\right\}$
$\approx 0.$
Lemma A.6.
Denoting the contexts from the source domain ${\mathcal{S}}$ as $\{{\bf x}_{t,a}^{\mathcal{S}}\}_{i\in[N],a\in[K]}$, the associated ground-truth action as $\{a_{i}^{*}\}_{i=1}^{N}$, we have the following inequality:
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\leq\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
where $\kappa_{i}^{\mathcal{S}}$ is defined in Lemma A.5.
Proof.
By definition, we have
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$
$=$ $\sum^{N}_{i=1}\left[\mathbbm{1}{[a_{i}^{*}=\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right]$
$\leq$ $\sum^{N}_{i=1}\left[\mathbbm{1}{[a_{i}^{*}=\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-1\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-0\right|\right)\right]$
$\leq$ $\sum^{N}_{i=1}\left[\mathbbm{1}{[a_{i}^{*}=\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-1\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right]$
$\leq$ $\sum^{N}_{i=1}\left[\mathbbm{1}{[a_{i}^{*}=\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-1\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-0\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right]$
[…]
$\equiv$ $\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
@@ A.2 Proof for the Target Regret Bound
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{2\cdot\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}}$
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$ (16)
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are the source regret, source-domain error, and ${\mathcal{H}\Delta\mathcal{H}}$ divergence defined in Definition 3.4,  Definition 3.3, and Definition 3.6, respectively. $\psi$ is a constant independent to the problem and and $C$ is a constant which can be ignored.
@@
Please check the full proof in the next page.
@@
$\mathit{R}_{\mathcal{T}}$ $=$ $\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)$
$=$ $\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)$
$\leq$ $\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)$
@@
$\leq$ ${\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\epsilon_{\mathcal{T}}(h)}+\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\right|\right)$
@@
$\leq$ ${\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\epsilon_{\mathcal{T}}(h)}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\left(\left|\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\right|\right)}$